Solve. Ln(2x^2+0.5)=0
Answer by stanbon(75887)

Solve. Ln(2x^2+0.5)=0
2x^2+0.5 = e^0
2x^2 + 0.5 = 1
2x^2 = 1/2
x^2 = 1/4
x = 1/2 or x = -1/2
===========================
Cheers,
Stan H.